OpenStudy (anonymous):

Solve the following system of equations. 3x - 2y = 6 6x - 4y = 14

OpenStudy (texaschic101):

3x - 2y = 6 --> (-2)3x - 2y = 6 6x - 4y = 14 ------------ -6x + 4y = - 12 (result of multiplying by -2) 6x - 4y = 14 ------------ 0 + 0 = 2 (incorrect) This system of equations has NO SOLUTION